A bag contains 6 marbles Three red marbles two blue marbles and one white marble If two marbles are selected at random determine the probability that both are red?
Answer by stanbon(48568)

A bag contains 6 marbles Three red marbles two blue marbles and one white marble If two marbles are selected at random determine the probability that both are red?
-----------------------
ways to get 2 red: 3C2 = 3
---------------
ways to get 2 of any color: 6C2 = 15
----------------------------
P(both red) = 3/15 = 1/5
===============================
Cheers,
Stan H.